OpenStudy (anonymous):

using the midpoint rule , approximate the integral: integral sign b=pi, a=0 sec(x/3)dx , n=6 Thanx

OpenStudy (anonymous):

if you carefully divide your interval, it should be \[ x_i=\frac{i\pi}{6} \] for \[ i=0,1,2,3,4,5,6\] Hence, the integral would become \[ \int_0^{\pi} \sec\left(\frac{x}{3}\right)dx =\frac{\pi}{12}\left(\sum_{i=0}^5\sec\left(\frac{x_i+x_{i+1}}{2}\right)\right)\]

OpenStudy (anonymous):

sorry, there is a typo, the thing in side sec on the right hand side should be divided by 6, not by 2.

OpenStudy (anonymous):

but don't we need to find delta x first ?

OpenStudy (anonymous):

delta x is \[ \frac{\pi}{12} \]

OpenStudy (anonymous):

so let me see if I got this right .....pi-0/6= pi/ 12 ?? shouldn't it be pi/6?

OpenStudy (anonymous):

Hang on, you are right, it should be pi/6, excuse me, I am on the rush

OpenStudy (anonymous):

ok so the xi=xi+1 should be divided by 2 not 6 right ?

OpenStudy (anonymous):

no, it is 6, indeed, you have totally 7 discrete points \[ x_0=0, x_1=\frac{\pi}{6}, x_2=\frac{2\pi}{6}, x_3=\frac{3\pi}{6},x_4=\frac{4\pi}{6},x_5=\frac{5\pi}{6},x_6=\pi \] and obviously, these 7 points will make up 6 intervals.

OpenStudy (anonymous):

so I can just go ahead and plug in those points into the original function right ? or I can add them up and divide by 6?

OpenStudy (anonymous):

yup, use the formula for the midpoint as I have given you, but it should be divided by 6, not by 2